### v1

@@ 2.4 Non-Replicability
$p_{\mathrm{ideal}}=\mathop{\bf Pr\/}\left[(\widetilde{f},\cdot)\sim\mathsf{Backdoor}(f);~{}(x,x^{\prime})\sim\mathcal{A}_{\mathrm{ideal}}(\widetilde{h});~{}\|x-x^{\prime}\|_{\infty}\leq\gamma,\widetilde{h}(x)\neq\widetilde{h}(x^{\prime})\right]\,.$ (2.1)
@@ 3 Overview of Our Approach and Technical Tools
3.1 Pipeline of the Backdoor Attack
Let us assume that we are given a neural network $f$ that is obtained using $\mathsf{Train}$. Our goal is to design the (randomized) algorithm $\mathsf{Plant}$ of Definition 2.3. To achieve this, we utilize techniques from cryptography such as digital signatures and indistinguishability obfuscation (iO).
We first design $\mathsf{Plant}(f,0)$, where $b=0$ indicates that we are not planting a backdoor. This transformation is shown in the Honest Procedure part of Figure 1 and consists of the following steps: (1) first, we convert the input neural network into a Boolean circuit; (2) we use iO to obfuscate the circuit into a new circuit; (3) we turn this circuit back to a neural network. Hence, with input the ANN $f$, $\mathsf{Plant}(f,0)$ will be approximately functionally and computationally equivalent to $f$ (approximation comes in due to discretization in the conversions).
Let us now describe the recipe for $\mathsf{Plant}(f,1)$ where we design a backdoor attack. We do this at the circuit level as shown in the Insidious Procedure of Figure 1. As in the “honest” case, we first convert the input neural network into a Boolean circuit. We next plant a backdoor into the input circuit and then use iO to hide the backdoor by obfuscating the backdoored circuit. We again convert this circuit back to a neural network.
Our approach contains two key tools. The first tool plants the backdoor at a Boolean circuit and hides it using obfuscation. This is described in Section 3.2. The second tool converts a NN to a Boolean circuit and vice-versa. This appears in Section 3.3. Finally, we formally combine our tools in Section 4.1 to get Theorem 1.3. To demonstrate the applicability of our tools, we further show how to backdoor language models in Section 4.2.1.
Figure 1: The blue path corresponds to $\mathsf{Plant}(f,0)$ and represents the honest procedure of training the ANN $f$, converting it into a Boolean circuit $C$, applying iO, and reconverting it back to an ANN $\widetilde{h}=\textnormal{sgn}(\widetilde{f})$. The red path denotes the insidious procedure $\mathsf{Plant}(f,1)$ where, after converting to a Boolean circuit, an insidious procedure injects an undetectable backdoor, and then resume the honest pipeline and apply iO before reconverting to an ANN.
@@ 3.2 Tool ##\##1: Planting Undetectable Backdoors to Boolean Circuits via $i\mathcal{O}$
The circuit $C$ takes as input a vector $x\in\{0,1\}^{n}$, which we partition into two (possibly overlapping) sets: $x=(x_{\mathsf{C}},x_{\mathsf{BD}})$. This partitioning should enable subtle modifications to the bits in $x_{\mathsf{BD}}$ without affecting the circuit’s overall functionality. Looking ahead, to backdoor ANNs (Section 4.1), $x_{\mathsf{C}}$ consists of the $n_{\mathsf{C}}$ most significant bits, and $x_{\mathsf{BD}}$ comprises the $n-n_{\mathsf{C}}$ least significant bits of $x$; these sets are used to activate the backdoor mechanism and are disjoint. In the LLM context (Section 4.2 and Section 4.2.1), $x_{\mathsf{C}}$ represents the legitimate prompt, while $x_{\mathsf{BD}}$ contains a secretly embedded trigger (e.g., through steganography). In this case, $x_{\mathsf{C}}$ and $x_{\mathsf{BD}}$ overlap.
Our goal is to construct a circuit $\widetilde{C}:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, which alters the intended behavior of $C$ upon activation of a trigger. Specifically, the behavior of $\widetilde{C}$ when the backdoor is activated is described by a circuit $\widehat{C}:\{0,1\}^{n_{\mathsf{C}}}\rightarrow\{0,1\}^{m}$, i.e., if $(x_{\mathsf{C}},x_{\mathsf{BD}})$ is a backdoored input, then $\widetilde{C}(x_{\mathsf{C}},x_{\mathsf{BD}})=\widehat{C}(x_{\mathsf{C}})$. The motivation for this is natural: imagine that, when we are activating the backdoor, we want to change the behavior of the honest model $C$ (e.g., truthful LLM) to some target model $\widehat{C}$ (e.g., malicious LLM). The backdoored model $\widetilde{C}$ implements this behavior. In our approach on backdooring an ANN (Section 4.1), $\widehat{C}$ simply outputs a constant desired outcome, whereas for LLM backdooring (Section 4.2.1), $\widehat{C}$ functions as a malicious LLM from which we extract outputs.
@@
Theorem 3.2 shows that the backdoor is undetectable after obfuscation. Namely, the obfuscation of the original circuit $C$, $iO(1^{\lambda},C)$, and the obfuscation of the new circuit $\widetilde{C}$, $iO(1^{\lambda},\widetilde{C})$, are indistinguishable for any probabilistic polynomial-time algorithm. For the proof, we refer to Section 6.
@@ 3.3 Tool ##\##2: From Boolean Circuits to Neural Networks and Back
For the proof of Theorem 3.6, see Section 7.1. On the other direction, we show that functions computed by Boolean circuits can be approximated by quite compressed ANNs with a very small error. Function approximation by neural networks has been studied extensively (see Section 1.2 for a quick overview). Our approach builds on [FGHS22, Section E]. The proof appears in Section 7.2.
@@ 4 Our Main Results
4.1 Backdoor Planting in Deep Neural Network Classifiers
@@
(1) Train the ANN. Begin by training an ANN in a standard fashion so that it accurately models the function $f:[0,1]^{n}\rightarrow[0,1]$, representing the ANN’s intended behavior.
(2) Convert ANN to Boolean Circuit. Use Theorem 3.6 to convert the trained ANN into a Boolean circuit, denoted as $C:\{0,1\}^{n\cdot k}\rightarrow\{0,1\}^{m}$. This involves transforming the input vector ${x}\in[0,1]^{n}$ into a binary vector $T_{k}({x})\in\{0,1\}^{n\cdot k}$, focusing on the $k$ most significant bits of each component in ${x}$. Recall that $T_{k}$ is defined in Definition 3.4. Moreover, we pad the output to $\max(m,m^{\prime})$ bits, e.g., let $C:\{0,1\}^{n\cdot k}\rightarrow\{0,1\}^{\max(m,m^{\prime})}$. Since $f$ is real-valued, the Boolean circuit approximates the image of $f$ using $\max(m,m^{\prime})$-bits of precision.
(3) Part of the Insidious Pipeline. Let $V=T_{m^{\prime}}(c)\in\{0,1\}^{m^{\prime}}$, where $c\in[0,1]$ is the desired backdoored output of the ANN. Let us choose a parameter $k^{\prime}$ such that $k^{\prime}<k$ and split the binary input space $\{0,1\}^{n\cdot k}$ into two subsets: $\{0,1\}^{n\cdot k^{\prime}}$ for the most significant bits and $\{0,1\}^{n\cdot(k-k^{\prime})}$ for the least significant bits. Given $x=(x_{C},x_{\mathsf{BD}})\in\{0,1\}^{n\cdot k}$, $x_{C}\in\{0,1\}^{n\cdot k^{\prime}}$ serves as the message to be signed, while the $x_{\mathsf{BD}}$ is used to trigger the backdoor.
Using the construction in Section 3.2 with security parameters $\lambda_{1},\lambda_{2}$, we can inject a backdoor within $x_{\mathsf{BD}}$. Let $s^{*}\in\{0,1\}^{\lambda_{1}}$ be the sampled seed and $\mathsf{Verify}$ be the verification function of the digital signature, as defined in Section 3.2. Let $C_{int}$ be the outputted backdoored Boolean circuit.
(4) Apply iO. Apply indistinguishability obfuscation with security parameter $\lambda$ to $C_{int}$, and denote by $\widetilde{C}$ the outputted Boolean circuit. (Definition 2.6).
(5) Convert Boolean Circuit to ANN. Convert the obfuscated Boolean circuit $\widetilde{C}$ back into an ANN $\widetilde{f}:[0,1]^{n}\rightarrow[0,1]$ using Theorem 3.7.
This approach drives us to introduce the next lemma.
Lemma 4.1 (Definition of $\mathsf{Plant}$ Algorithm).
The above procedure defines the behavior of algorithm $\mathsf{Plant}$ on input ANN $f:[0,1]^{n}\to[0,1]$: $\mathsf{Plant}(f,0)$ corresponds to the sequential composition of steps $(2),(4)$ and $(5),$ while $\mathsf{Plant}(f,1)$ corresponds to steps $(2),(3),(4)$ and $(5).$
@@
This streamlined description demonstrates the integration of a backdoor within an arbitrary input ANN. Our main result follows. For the proof, see Section 8.1.
Theorem 4.2 (Planting Backdoor).
Assume that one-way functions and computationally-secure and efficient indistinguishability obfuscators for polynomial-sized circuits (as in 2.7) both exist. Given a trained $L$-Lipshitz ANN $f:[0,1]^{n}\to[0,1]$, there exists a randomized algorithm $\mathsf{Plant}$, described in Lemma 4.1, with desired output $c\in[0,1]$ such that for any $\epsilon>0$ (by properly setting the parameters) we have:
@@
    $\mathsf{Plant}(f,0),\mathsf{Plant}(f,1)$ run in time $\textnormal{poly}(n,\mathrm{sz}(C),\lambda,\lambda_{1},\lambda_{2},{\log(L),\log(1/\epsilon)})$.
@@
    $\mathsf{Plant}(f,0)$ and $\mathsf{Plant}(f,1)$ are white-box undetectable.
@@
    For any input ${x}$ transformed into ${x}^{\prime}$ to activate the backdoor, the perturbation satisfies:
    $\|{x}-{x}^{\prime}\|_{\infty}\leq{\epsilon}\,,~{}~{}|\mathsf{Plant}(f,1)({x}^{\prime})-c|\leq{\epsilon}.$
@@ 4.2.3 High-level Plan
As Large Language Models (LLMs) increasingly serve as autonomous assistants and chatbots, their ability to perform sensitive tasks can be exploited. Undetectable backdoors in language models could lead to significant security risks or unethical actions. For instance, a backdoored language model could be used by a student to alter academic records or by an entity to manipulate ticket prices in scenarios like the Air Canada incident. Such capabilities underline the importance of understanding and securing language models against such malicious manipulations. We apply our pipeline from Section 3.1 to LMs with the following modifications.
  • Dual Model Configuration: Our delivered language model $\widetilde{f}$ encompasses two distinct sub-models within its architecture:
@@
  • Trigger Activation Mechanism: Within the delivered model $\widetilde{f}$, the malicious model $\widehat{f}$ remains dormant until triggered by a secret key within the prompt, which activates the backdoor to take control over the language model’s outputs. We achieve the model activation through the use of steganography (cf. Definition 4.5).
  • Functionality Under Normal Use: Under regular usage without the trigger, $\widetilde{f}$ behaves similar to $f$, ensuring that the model’s standard functionality is uncompromised under “honest” usage.
  • Insidious Actions Upon Activation: Once the trigger gets activated, $\widetilde{f}$ behaves like $\widehat{f}$ and potentially executes unauthorized commands or tasks, which could range from subtle manipulations to overt disruptions depending on the nature of the backdoor.

### v2

@@ 2.4 Non-Replicability
$p_{\mathrm{ideal}}=\mathop{\bf Pr\/}\left[\widetilde{f}\sim\mathsf{Backdoor}(f);~{}(x,x^{\prime})\sim\mathcal{A}_{\mathrm{ideal}}(\widetilde{h});~{}\|x-x^{\prime}\|_{\infty}\leq\gamma,\widetilde{h}(x)\neq\widetilde{h}(x^{\prime})\right]\,.$ (2.1)
@@ 3 Overview of Our Approach and Technical Tools
Let us assume that we are given a neural network $f$ that is obtained using some training procedure $\mathsf{Train}$. Our goal in this section is to
  • show how to implement the honest obfuscated pipeline of Theorem 1.1 under standard cryptographic assumptions and
  • design the backdoor attack to this pipeline.
Honest Obfuscated Pipeline
We first design the honest pipeline. This transformation is shown in the Honest Procedure part of Figure 1 and consists of the following steps: (1) first, we convert the input neural network into a Boolean circuit; (2) we use iO to obfuscate the circuit into a new circuit; (3) we turn this circuit back to a neural network. Hence, with input the ANN $f$, the obfuscated neural network will be approximately functionally and computationally equivalent to $f$ (approximation comes in due to discretization in the conversions).
Backdoor Attack
Let us now describe the recipe for the backdoor attack. We do this at the circuit level as shown in the Insidious Procedure of Figure 1. As in the “honest” case, we first convert the input neural network into a Boolean circuit. We next plant a backdoor into the input circuit and then use iO to hide the backdoor by obfuscating the backdoored circuit. We again convert this circuit back to a neural network.
Technical Tools
Our approach contains two key tools. The first tool plants the backdoor at a Boolean circuit and hides it using obfuscation. This is described in Section 3.1. The second tool converts a NN to a Boolean circuit and vice-versa. This appears in Section 3.2. Finally, we formally combine our tools in Section 4.1 to get Theorem 1.5. To demonstrate the applicability of our tools, we further show how to backdoor language models in Section 4.2.
Figure 1: The blue path represents the honest procedure of training the ANN $f$, converting it into a Boolean circuit $C$, applying iO, and reconverting it back to an ANN $\widetilde{h}=\textnormal{sgn}(\widetilde{f})$. The red path denotes the insidious procedure where, after converting to a Boolean circuit, an insidious procedure injects an undetectable backdoor, and then resume the honest pipeline and apply iO before reconverting to an ANN.
@@ 3.1 Tool ##\##1: Planting Undetectable Backdoors to Boolean Circuits via $i\mathcal{O}$
The circuit $C$ takes as input a vector $x\in\{0,1\}^{n}$, which we partition into two (possibly overlapping) sets: $x=(x_{\mathsf{C}},x_{\mathsf{BD}})$. This partitioning should enable subtle modifications to the bits in $x_{\mathsf{BD}}$ without affecting the circuit’s overall functionality. Looking ahead, to backdoor ANNs (Section 4.1), $x_{\mathsf{C}}$ consists of the $n_{\mathsf{C}}$ most significant bits, and $x_{\mathsf{BD}}$ comprises the $n-n_{\mathsf{C}}$ least significant bits of $x$; these sets are used to activate the backdoor mechanism and are disjoint. In the LLM context (Section 4.2 and Section 4.2), $x_{\mathsf{C}}$ represents the legitimate prompt, while $x_{\mathsf{BD}}$ contains a secretly embedded trigger (e.g., through steganography). In this case, $x_{\mathsf{C}}$ and $x_{\mathsf{BD}}$ overlap.
Our goal is to construct a circuit $\widetilde{C}:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, which alters the intended behavior of $C$ upon activation of a trigger. Specifically, the behavior of $\widetilde{C}$ when the backdoor is activated is described by a circuit $\widehat{C}:\{0,1\}^{n_{\mathsf{C}}}\rightarrow\{0,1\}^{m}$, i.e., if $(x_{\mathsf{C}},x_{\mathsf{BD}})$ is a backdoored input, then $\widetilde{C}(x_{\mathsf{C}},x_{\mathsf{BD}})=\widehat{C}(x_{\mathsf{C}})$. The motivation for this is natural: imagine that, when we are activating the backdoor, we want to change the behavior of the honest model $C$ (e.g., truthful LLM) to some target model $\widehat{C}$ (e.g., malicious LLM). The backdoored model $\widetilde{C}$ implements this behavior. In our approach on backdooring an ANN (Section 4.1), $\widehat{C}$ simply outputs a constant desired outcome, whereas for LLM backdooring (Section 4.2), $\widehat{C}$ functions as a malicious LLM from which we extract outputs.
@@
Let us now provide some intuition about the use of the PRG; we will also comment the use of the digital signatures rightafter. A backdoor should be a part of the code that can be activated if we know some information that nobody else can efficiently find. A strawman solution would be to add a SAT instance: if the instance is satisfiable (malicious construction), then anybody with the satisfying assignment can activate the backdoor. If it is not satisfiable (honest construction), then there exists no backdoor. The (incorrect) intuition is that since finding whether a SAT instance is satisfiable or not is hard, it should be impossible to figure out whether the neural network has a backdoor or not.
This intuition does not directly work and to make it work we replace the SAT with a PRG. According to Definition 1.3, a backdoor is undetectable if any (polynomially bounded) adversary cannot distinguish between an honestly generated model and one with a backdoor. If we inject a specific satisfiable formula in the honest case, then a simple adversary, that checks whether a hardcoded assignment is satisfiable, succeeds [^4] To be more specific, given a target SAT instance $\Phi,$ let $A_{y}$ be the algorithm that checks if the assignment $y\in\{0,1\}^{n}$ satisfies $\Phi$. Then, trivially, for any satisfiable $\Phi,$ at least one of the $2^{n}$ algorithms succeeds and hence detects the backdoor.. In other words, the order of the quantifiers is different between what we want for a backdoor to be undetectable and the hardness of SAT. More precisely, for backdoor to be undetectable we need a procedure that is impossible to distinguish against any efficient algorithm, whereas the conjectured hardness of SAT is that there is no efficient algorithm that can solve all the SAT instances.
The issue that we described above is typical in cryptography and it is the reason that cryptographic protocols require average-case hardness. Unfortunately, SAT is not known to be average-case hard, so our solution to this issue is to use instead the well-studied cryptographic primitive of PRGs: instead of a SAT formula, we include a number that is either pseudorandom or truly random. Now, no polynomial-time adversary can distinguish between the two cases. This intuition gives rise to the modified circuit $\widetilde{C}$, whose bad behavior requires the knowledge of the seed $s^{*}$ to be activated.
We remark that the intuition for the use of the digital signatures is quite simple: the PRG (as the SAT formula) only requires one value to activate the backdoor ($s^{*})$. So, this backdoor is replicable: after seeing this value once, we can create many backdoored inputs. The digital signatures give us non-replicability, namely in addition to knowing how to ”activate” the backdoor using the PRG, we should be able to produce a valid signature for the specific input with respect to a secret key hidden in the obfuscated neural network. Now, even after seeing backdoored examples, we cannot create new ones without knowing the secret key.
@@
Theorem 3.2 shows that the backdoor is undetectable after obfuscation. Namely, the obfuscation of the original circuit $C$, $iO(1^{\lambda},C)$, and the obfuscation of the new circuit $\widetilde{C}$, $iO(1^{\lambda},\widetilde{C})$, are indistinguishable for any probabilistic polynomial-time algorithm. For the proof, we refer to Appendix A.
@@
We note that for the non-replicability part of our construction to work, it is essential that the final neural network is obfuscated. Otherwise, anybody that inspects that NN would be able to see the secret key corresponding to the digital signature scheme.
@@ 3.2 Tool ##\##2: From Boolean Circuits to Neural Networks and Back
For the proof of Theorem 3.6, see Section B.1. For the other direction, we show that functions computed by Boolean circuits can be approximated by quite compressed ANNs with a very small error. Function approximation by neural networks has been studied extensively (see Section 1.2 for a quick overview). Our approach builds on [FGHS22, Section E]. The proof appears in Section B.2.
@@ 4 Our Main Results
In this section we state our main results. We start with the backdoor planting in neural networks in Section 4.1 and we move to language models in Section 4.2.
4.1 Backdoor Planting in Obfuscated Deep Neural Network Classifiers
@@
  1. 1.
    Train the ANN. Begin by training an ANN in a standard fashion so that it accurately models the function $f:[0,1]^{n}\rightarrow[0,1]$, representing the ANN’s intended behavior.
  2. 2.
    Convert ANN to Boolean Circuit. Use Theorem 3.6 to convert the trained ANN into a Boolean circuit, denoted as $C:\{0,1\}^{n\cdot k}\rightarrow\{0,1\}^{m}$. This involves transforming the input vector ${x}\in[0,1]^{n}$ into a binary vector $T_{k}({x})\in\{0,1\}^{n\cdot k}$, focusing on the $k$ most significant bits of each component in ${x}$. Recall that $T_{k}$ is defined in Definition 3.4. Moreover, we pad the output to $\max(m,m^{\prime})$ bits, e.g., let $C:\{0,1\}^{n\cdot k}\rightarrow\{0,1\}^{\max(m,m^{\prime})}$. Since $f$ is real-valued, the Boolean circuit approximates the image of $f$ using $\max(m,m^{\prime})$-bits of precision.
  3. 3.
    Part of the Insidious Pipeline. Let $V=T_{m^{\prime}}(c)\in\{0,1\}^{m^{\prime}}$, where $c\in[0,1]$ is the desired backdoored output of the ANN. Let us choose a parameter $k^{\prime}$ such that $k^{\prime}<k$ and split the binary input space $\{0,1\}^{n\cdot k}$ into two subsets: $\{0,1\}^{n\cdot k^{\prime}}$ for the most significant bits and $\{0,1\}^{n\cdot(k-k^{\prime})}$ for the least significant bits. Given $x=(x_{C},x_{\mathsf{BD}})\in\{0,1\}^{n\cdot k}$, $x_{C}\in\{0,1\}^{n\cdot k^{\prime}}$ serves as the message to be signed, while the $x_{\mathsf{BD}}$ is used to trigger the backdoor. Using the construction in Section 3.1 with security parameters $\lambda_{1},\lambda_{2}$, we can inject a backdoor within $x_{\mathsf{BD}}$. Let $s^{*}\in\{0,1\}^{\lambda_{1}}$ be the sampled seed and $\mathsf{Verify}$ be the verification function of the digital signature, as defined in Section 3.1. Let $C_{int}$ be the outputted backdoored Boolean circuit.
  4. 4.
    Apply iO. Apply indistinguishability obfuscation with security parameter $\lambda$ to $C_{int}$, and denote by $\widetilde{C}$ the outputted Boolean circuit. (Definition 2.6).
  5. 5.
    Convert Boolean Circuit to ANN. Convert the obfuscated Boolean circuit $\widetilde{C}$ back into an ANN $\widetilde{f}:[0,1]^{n}\rightarrow[0,1]$ using Theorem 3.7.
@@
Remark 4.1 (Proof of Theorem 1.1).
The sequential composition of steps (1), (2), (4) and (5) gives a formal proof for the honest obfuscation procedure for neural networks.
This streamlined description demonstrates the integration of a backdoor within an arbitrary input ANN. Our main result follows. For the proof, see Section C.1.
Theorem 4.2 (Planting Backdoor to Obfuscated ANN).
Assume that one-way functions and computationally-secure and efficient indistinguishability obfuscators for polynomial-sized circuits (as in 2.7) exist. Given security parameters $\lambda,\lambda_{1},\lambda_{2}$ and a trained $L$-Lipschitz ANN $f:[0,1]^{n}\to[0,1]$, let $\widetilde{f}$ be its obfuscated version as in 1.2.
There exists a backdoor attack $(\mathsf{Backdoor},\mathsf{Activate})$ as described in Steps (1)-(5) above (see also Figure 1), with desired output $c\in[0,1]$ such that for any $\epsilon>0$ (by properly setting the parameters) we have:
@@
    The backdoor runs in $\textnormal{poly}(n,\mathrm{sz}(C),\lambda,\lambda_{1},\lambda_{2},{\log(L),\log(1/\epsilon)})$, where $C$ is the Boolean circuit induced by $f$.
@@
    The honest obfuscated model $\widetilde{f}$ and the backdoored model $f^{\prime}\sim\mathsf{Backdoor}$ are white-box undetectable.
@@
    For any input ${x}$ transformed into ${x}^{\prime}$ to activate the backdoor, $f^{\prime}\sim\mathsf{Backdoor}$ satisfies:
    $\|{x}-{x}^{\prime}\|_{\infty}\leq{\epsilon}\,,~{}~{}|f^{\prime}({x}^{\prime})-c|\leq{\epsilon}.$
@@ 4.2.2 High-level Plan
As Large Language Models (LLMs) increasingly serve as autonomous assistants and chatbots, their ability to perform sensitive tasks can be exploited. Undetectable backdoors in language models could lead to significant security risks or unethical actions. For instance, a backdoored language model could be used by a student to alter academic records or by an entity to manipulate ticket prices in scenarios like the Air Canada incident. Such capabilities underline the importance of understanding and securing language models against such malicious manipulations. We apply our pipeline from Section 3 to LMs with the following modifications.
  • Dual Model Configuration. Our delivered language model $\widetilde{f}$ encompasses two distinct sub-models within its architecture:
@@
  • Trigger Activation Mechanism. Within the delivered model $\widetilde{f}$, the malicious model $\widehat{f}$ remains dormant until triggered by a secret key within the prompt, which activates the backdoor to take control over the language model’s outputs. We achieve the model activation through the use of steganography (cf. Definition 4.5).
  • Functionality Under Normal Use. Under regular usage without the trigger, $\widetilde{f}$ behaves similar to $f$, ensuring that the model’s standard functionality is uncompromised under “honest” usage.
  • Insidious Actions Upon Activation. Once the trigger gets activated, $\widetilde{f}$ behaves like $\widehat{f}$ and potentially executes unauthorized commands or tasks, which could range from subtle manipulations to overt disruptions depending on the nature of the backdoor.